# Measuring What Makes You Unique: Difference-Aware User Modeling for Enhancing LLM Personalization

### v1

@@ Abstract
Yilun Qiu1††thanks: Equal contribution, Xiaoyan Zhao211footnotemark: 1, Yang Zhang1††thanks: Corresponding Author, Yimeng Bai3, Wenjie Wang3, Hong Cheng2, Fuli Feng3, Tat-Seng Chua1, 1National University of Singapore 2The Chinese University of Hong Kong 3University of Science and Technology of China qiuyilun@u.nus.edu, xzhao@se.cuhk.edu.hk, zyang1580@gmail.com, baiyimeng@mail.ustc.edu.cn, wenjiewang96@gmail.com, hcheng@se.cuhk.edu.hk, fulifeng93@gmail.com, dcscts@nus.edu.sg
@@ 1 Introduction
With continuous efforts, Large Language Models (LLMs) Achiam et al. (2023); Yang et al. (2024); Guo et al. (2025); Dubey et al. (2024) have demonstrated increasingly higher levels of intelligence, sparking an unprecedented enthusiasm for applying them to individuals’ daily lives to enhance personal thinking, planning, and overall life experiences Bao et al. (2023); Razafinirina et al. (2024); Li et al. (2024b); Deng et al. (2024); Zhang et al. (2025b). However, the “one-size-fits-all” paradigm of general LLMs becomes ineffective when serving individual users, as users have distinct preferences and would expect personalized generation from LLMs Christakopoulou et al. (2023). This motivates an exciting new direction: LLM personalization Zhang et al. (2024b) – adapting LLMs to generate text responses that align with each individual’s unique preferences.
LLM personalization fundamentally relies on incorporating user-specific information into LLMs to tailor their generation Kang et al. (2023); Li et al. (2023b). The “memorize-then-inject" framework offers a promising and economical solution – storing users’ historical data and then extracting key information to serve as instructional contexts for customizing LLM responses Salemi et al. (2024b). Depending on the extraction method employed, existing approaches can generally be categorized into: 1) Retrieval-based methods Mysore et al. (2024); Li et al. (2023a), which just retrieve the parts of the user’s history relevant to the current request for LLM use, and 2) Summarization-based methods Richardson et al. (2023); Zheng et al. (2024), which further condense the user history to distill preferences for LLM use.
@@
We validate our DPL method on a representative personalized text generation task, namely, review generation Peng et al. (2024). Extensive results demonstrate that DPL effectively enhances LLM personalization.
@@ 2 Preliminary
where denotes the generated result, denotes the historical data for , and represents the key information extraction process from . Notably, the extraction process usually also involves the utilization of an LLM. This method does not require LLM retraining and can potentially manage lifelong historical data in a memory manner Zhang et al. (2024a), making it a promising and cost-effective solution.
@@ 3.2.1 Selective User Comparison
To select representative users for comparison, the core consideration is that they could represent the total population, making the comparison still provide a global view. To achieve this, we consider clustering the users based on their text and then selecting the user in the cluster center for comparison. Specifically, when dealing with each key element in history for the target , we just consider the users having texts on , and then cluster these users into groups according to the texts corresponding to them, obtaining the most central users for each group. We then select the central users that do not belong to groups of the target user as the users for comparison. Formally,
@@ 3.2.2 Structured Difference Extraction
Extracting differences along these dimensions ensures meaningful and well-structured results for distinguishing task-relevant content. We guide the LLM to adhere to these dimensions for difference extraction using prompts. All prompts used in DPL are provided in Appendix F.
@@ 4 Experiments
| Datasets () | Books | Movies & TV | CDs & Vinyl | |||||||||
| Methods () | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT | R-1 | MET. | S-72B | S-GPT |
| Non-Perso | 0.2683 | 0.1512 | 0.5079 | 0.4360 | 0.2275 | 0.1304 | 0.4767 | 0.3242 | 0.2392 | 0.1336 | 0.4701 | 0.3409 |
| RAG | 0.3206 | 0.2173 | 0.6044 | 0.4700 | 0.2845 | 0.1897 | 0.5429 | 0.3447 | 0.2926 | 0.1876 | 0.5328 | 0.3701 |
| IntSum | 0.3174 | 0.2304 | 0.6177 | 0.4826 | 0.2846 | 0.1965 | 0.5728 | 0.3521 | 0.2963 | 0.1979 | 0.5509 | 0.3714 |
| LLM-TRSR | 0.3143 | 0.2294 | 0.6189 | 0.4770 | 0.2822 | 0.1955 | 0.5781 | 0.3571 | 0.2954 | 0.1979 | 0.5525 | 0.3719 |
| CICL | 0.3267 | 0.2370 | 0.6233 | 0.4978 | 0.2868 | 0.1971 | 0.5684 | 0.3618 | 0.2980 | 0.1968 | 0.5510 | 0.3677 |
| Persona-DB | 0.3279 | 0.2430 | 0.6259 | 0.4839 | 0.2836 | 0.1956 | 0.5728 | 0.3589 | 0.2936 | 0.1939 | 0.5481 | 0.3669 |
| DPL (ours) | 0.3326 | 0.2445 | 0.6278 | 0.5136 | 0.2934 | 0.2011 | 0.5795 | 0.3812 | 0.3061 | 0.2061 | 0.5684 | 0.3780 |
@@ 4.1 Experimental Setup
Datasets. We focus on review generation, a representative personalized text generation task. Therefore, we conduct experiments on the Amazon Reviews 2023 dataset Hou et al. (2024)111https://amazon-reviews-2023.github.io/, focusing on the categories of Books, Movies & TV, and CDs & Vinyl. This dataset aggregates user-item interactions from Amazon, encompassing user reviews (ratings, texts, etc.) and item metadata (titles, descriptions, etc.). We preprocess the data into a format suitable for personalized review generation tasks, with details provided in Appendix A.
Baselines. We compare DPL with the following methods, with more details provided in Appendix B. Note that all methods are implemented using the same template and input format, as outlined in Appendix F, to ensure fair comparisons.
@@
Conventional Evaluation. Following previous work on personalized text generation Kumar et al. (2024), we employ widely-adopted ROUGE-1 Lin (2004) and METEOR Banerjee and Lavie (2005) to measure lexical overlap between the generated reviews and ground-truth reviews, offering surface-level comparisons through n-gram matching and semantic alignment mechanisms Wang et al. (2023).
@@
Implementation Details. For our DPL, we use BM25 Robertson and Zaragoza (2009) as the retriever and Qwen2.5-14B-Instruct Yang et al. (2024) as the LLM backbone, which functions as the difference extractor, summarizer, and review generator. For selective user comparison and structured difference extraction, we set the number of retrieved reviews, i.e., in Equation (3), to 8. Additionally, we utilize gte-Qwen2-1.5B-Instruct Li et al. (2023c) to map user reviews into embeddings, which are subsequently clustered into 5 groups by default, equivalent to setting the number of selected representative users in Equation (4) to 4. More details are provided in Appendix D.
@@ 4.2 Main Results (RQ1)
DPL demonstrates superior personalization performance compared to the baseline methods across all metrics on three datasets. For instance, on the CDs & Vinyl dataset, it shows an average improvement of 2.84% over the best baseline metrics. This result underscores its enhanced personalized text generation capabilities, which can be attributed to its focus on extracting inter-user differences to improve LLM personalization through the selective comparison mechanism and difference extraction standard.
@@ 4.3 Abalation Studies (RQ2)
Random. This variant randomly selects users without considering any specific criteria.
@@
| Methods | R-1 | MET. | S-72B | S-GPT |
| DPL | 0.3326 | 0.2445 | 0.6278 | 0.5136 |
| Random | 0.3282 | 0.2392 | 0.6246 | 0.5041 |
| SimRank | 0.3301 | 0.2421 | 0.6233 | 0.5078 |
| None | 0.3212 | 0.2362 | 0.6221 | 0.4877 |
| WriOnly | 0.3322 | 0.2461 | 0.6120 | 0.4959 |
| EmoOnly | 0.3241 | 0.2379 | 0.6151 | 0.5016 |
| SemOnly | 0.3244 | 0.2382 | 0.6196 | 0.5066 |
Table 2 illustrates the comparison results, from which we draw the following observations:
@@
For the structured difference extraction, removing the parts of the complete extraction dimensions leads to a general performance decline. This finding suggests that each dimension plays a crucial role in capturing different facets of the differences, thereby ensuring meaningful and well-structured extraction outcomes for distinguishing task-relevant content.
@@ 4.4.1 Hyper-parameter Analysis
In our investigation, the hyperparameter , which denotes the number of selected representative users for comparison, plays a pivotal role in determining the effectiveness of DPL. We vary within the set to assess the impact of varying these parameters on performance. Additionally, we present the best baseline metrics for comparative analysis, with the results shown in Figure 2. Our findings indicate that, generally, increasing the number of representative users leads to slight to moderate improvements in performance, with the best results occurring at . This underscores that effective difference extraction depends on selecting an appropriate number of comparative representatives.
@@ 4.4.2 Impact of User Uniqueness
Following this classification, we compare the performance of DPL with the non-personalization method across the two groups. For simplicity, we report the absolute improvement of R-1 and S-72B, as the trends for other metrics are consistent. As illustrated in Figure 3, the unique group shows a significantly greater improvement than the non-unique group. This suggests that the more unique the user, the better we can capture insights into their differences relative to others, aligning with the philosophy that “differences make us unique”.
@@ 4.5 Case Study (RQ5)
Finally, we conduct a case study to gain a deeper understanding of DPL’s capacity for personalized review generation. Specifically, we select a representative item from the Movies & TV dataset, and compare the reviews generated by DPL with those produced by the representative baseline method, LLM-TRSR. The detailed reviews are presented in Figure 4, where we highlight consistent sections between the generated reviews and the user’s real review by distinct colors.
We observe that the review generated by DPL exhibits more consistent sections. Furthermore, it offers a more personalized perspective and concise content, whereas the review generated by LLM-TRSR adopts a more objective stance and contains more tedious information. The alignment of DPL with the user’s writing style and semantic preferences further supports its superiority. More case studies are provided in Appendix G.
@@ 5 Related Work
LLM personalization has been explored across various domains, underscoring its importance. While existing works such as PUMA Cai et al. (2024) and PMG Shen et al. (2024) advance personalization in different scenarios, personalized text generation remains the foundation of LLM personalization Zhang et al. (2024b). LaMP Salemi et al. (2024b) evaluates short-form personalized text, later extended by LongLaMP Kumar et al. (2024) to assess long-form text generation. Among these, personalized item review writing stands out as a key task Peng et al. (2024) since user reviews naturally reflect individual preferences Ni et al. (2019), serving as a critical scenario to evaluate a model’s ability to generate personalized content.
Recent LLM-based enhance personalization using user-specific content, which can be categorized into: 1) retrieval-based methods, where ROPG Salemi et al. (2024a) introduces a retrieval selection model to dynamically choose the best retriever and HYDRA Zhuang et al. (2024) employs a personalized reranker to prioritize useful information; and 2) summarization-based methods, where ONCE Liu et al. (2024b) generates user profiles by summarizing topics and regions of interest in their browsing history and PPlug Liu et al. (2024a) converts user history into personalized embeddings.
Besides, a recent study Persona-DB Sun et al. (2025) has explored using data from other users to enhance personalized text generation. However, this method mainly focuses on enriching relevant knowledge rather than capturing each user’s uniqueness, which is crucial for effective personalization. To leverage inter-user data, contrastive learning Gao and Das (2024) offers a promising way of using both positive and negative examples to better describe intents. However, its performance is contingent on the identification of meaningful negative examples, which are challenging to obtain in user-centered LLM personalization. This reliance limits the full leverage of inter-user information for LLM personalization. To the best of our ability, our proposed DPL addresses this limitation.
@@ 6 Conclusions
In our research, we utilize an open-source dataset, ensuring full compliance with the original dataset’s MIT license. Our research prioritizes data integrity, privacy, and ethical considerations to uphold the responsible and fair use of open-source resources.
@@ Appendix A Dataset Details
Data Curation. We perform a detailed data-cleaning process to create a high-quality benchmark. First, we ensure that each review includes a corresponding review_title, rating, and timestamp, with a minimum text length of 200 characters to meet the long-form criteria. For each item, we verify that it has an item_title and that its description falls within the 100 to 2000-character range to maintain high-quality content. Following this, we ensure that each product has at least four unique reviewers and retain only the most recent review from each reviewer for a given item. Additionally, we restrict the number of reviews per reviewer to between 18 and 500 to filter out users engaging in review manipulation. These settings are designed to align with our experiments.
Dataset Usage. In our experiments, we utilize only the test dataset for a direct evaluation of model performance. The details of model inputs are provided in Appendix F. Notably, the training and validation datasets are included to support related research.
@@
Appendix B Baselines Details

### v2

@@ Abstract
Yilun Qiu1††thanks: Equal contribution, Xiaoyan Zhao211footnotemark: 1, Yang Zhang1††thanks: Corresponding Author, Yimeng Bai3, Wenjie Wang3, Hong Cheng2, Fuli Feng3, Tat-Seng Chua1 1National University of Singapore 2The Chinese University of Hong Kong 3University of Science and Technology of China qiuyilun@u.nus.edu, xzhao@se.cuhk.edu.hk, zyang1580@gmail.com, baiyimeng@mail.ustc.edu.cn, wenjiewang96@gmail.com, hcheng@se.cuhk.edu.hk, fulifeng93@gmail.com, dcscts@nus.edu.sg
@@ 1 Introduction
With continuous efforts, Large Language Models (LLMs) Achiam et al. (2023); Yang et al. (2024); Dubey et al. (2024); Guo et al. (2025) have demonstrated increasingly higher levels of intelligence, sparking an unprecedented enthusiasm for applying them to individuals’ daily lives to enhance personal thinking, planning, and overall life experiences Bao et al. (2023); Li et al. (2024a); Deng et al. (2024); Zhang et al. (2025b). However, the “one-size-fits-all” paradigm of general LLMs becomes ineffective when serving individual users, as users have distinct preferences and would expect personalized generation from LLMs Christakopoulou et al. (2023). This motivates an exciting new direction: LLM personalization Zhang et al. (2024d); Xu et al. (2025b); Liu et al. (2025a); Chen et al. (2024a) – adapting LLMs to generate text responses that align with each individual’s unique preferences.
LLM personalization fundamentally relies on incorporating user-specific information into LLMs to tailor their generation Chen et al. (2025); Zhao et al. (2025d). The “memorize-then-inject" framework offers a promising and economical solution – storing users’ historical data and then extracting key information to serve as instructional contexts for customizing LLM responses Salemi et al. (2024b); Wang et al. (2024); Wu et al. (2025). Depending on the extraction method employed, existing approaches can generally be categorized into: 1) Retrieval-based methods Li et al. (2023a); Mysore et al. (2024), which just retrieve the parts of the user’s history relevant to the current request for LLM use, and 2) Summarization-based methods Richardson et al. (2023); Zheng et al. (2024), which further condense the user history to distill preferences for LLM use.
@@
We validate our DPL method on a representative personalized text generation task, namely, review generation Ni et al. (2019); Peng et al. (2024); Au et al. (2025). Extensive results demonstrate that DPL effectively enhances LLM personalization.
@@ 2 Preliminary
where denotes the generated result, denotes the historical data for , and represents the key information extraction process from . Notably, the extraction process usually also involves the utilization of an LLM. This method does not require LLM retraining and can potentially manage lifelong historical data in a memory manner Zhang et al. (2024c), making it a promising and cost-effective solution.
@@ 3.2.1 Selective User Comparison
To select representative users for comparison, the core consideration is that they could represent the total population, making the comparison still provide a global view. To achieve this, we consider clustering the users based on their text and then selecting the user in the cluster center for comparison. Specifically, when dealing with each key element in history for the target , we just consider the users having texts on , and then cluster these users into groups using the K-means algorithm MacQueen (1967) according to the texts corresponding to them, obtaining the most central users for each group. We then select the central users that do not belong to groups of the target user as the users for comparison. Formally,
@@ 3.2.2 Structured Difference Extraction
Extracting differences along these dimensions ensures meaningful and well-structured results for distinguishing task-relevant content. We guide the LLM to adhere to these dimensions for difference extraction using prompts. All prompts used in DPL are provided in Appendix G.
@@ 4.1 Experimental Setup
Datasets. We focus on review generation, a representative personalized text generation task. So we conduct experiments on the Amazon Reviews 2023 dataset Hou et al. (2024)111https://amazon-reviews-2023.github.io/, focusing on the categories of Movies & TV, CDs & Vinyl, and Books. This dataset aggregates user-item interactions from Amazon, encompassing user reviews (ratings, texts, etc.) and item metadata (titles, descriptions, etc.). We preprocess the data into a format suitable for personalized review generation tasks, with details provided in Appendix A. Our processed datasets are publicly available on Huggingface222https://huggingface.co/datasets/SnowCharmQ/DPL-main & https://huggingface.co/datasets/SnowCharmQ/DPL-meta.
Baselines. We compare DPL with the following methods, with more details provided in Appendix B. Note that all methods are implemented using the same template and input format, as outlined in Appendix G, to ensure fair comparisons.
@@
Conventional Evaluation. Following previous work on personalized text generation Kumar et al. (2024), we employ widely-adopted ROUGE-1 Lin (2004), ROUGE-L, METEOR Banerjee and Lavie (2005), and BLEU Papineni et al. (2002) to measure lexical overlap between the generated reviews and ground-truth reviews, offering surface-level comparisons through n-gram matching and semantic alignment mechanisms Wang et al. (2023).
@@
Datasets () | Movies & TV | CDs & Vinyl | Books | ||||||
Methods () | ROUGE-1 | METEOR | S-72B | ROUGE-1 | METEOR | S-72B | ROUGE-1 | METEOR | S-72B |
Non-Perso | |||||||||
RAG | |||||||||
IntSum | |||||||||
LLM-TRSR | |||||||||
CICL | |||||||||
Persona-DB | |||||||||
DPL (ours) | |||||||||
Datasets () | Movies & TV | CDs & Vinyl | Books | ||||||
Methods () | ROUGE-L | BLEU | S-GPT | ROUGE-L | BLEU | S-GPT | ROUGE-L | BLEU | S-GPT |
Non-Perso | |||||||||
RAG | |||||||||
IntSum | |||||||||
LLM-TRSR | |||||||||
CICL | |||||||||
Persona-DB | |||||||||
DPL |
Implementation Details. For our DPL, we use the Qwen2.5-14B-Instruct Yang et al. (2024) model as the LLM backbone, which functions as the difference extractor, summarizer, and review generator. Following the setting of LongLaMP Kumar et al. (2024), we retrieve user reviews as auxiliary signals using BM25, where the retrieval query is formed by concatenating a short review title with the item’s description. For selective user comparison and structured difference extraction, we set the number of retrieved reviews, i.e., in Equation (3), to 8. Additionally, we utilize gte-Qwen2-1.5B-Instruct Li et al. (2023b) to map user reviews into embeddings, which are subsequently clustered into 5 groups by default, equivalent to setting the number of selected representative users in Equation (4) to 4. To assess the reliability of our generated results, we conduct inference using different temperature values (0.2, 0.4, 0.6, 0.8, and 1.0) for each method, and report the mean and standard deviation of the evaluation metrics. We also perform t-tests to assess statistical significance. More details are provided in Appendix D.
@@ 4.2 Main Results (RQ1)
DPL demonstrates consistently superior personalization performance across all three datasets and all evaluation metrics. For instance, on the CDs & Vinyl category, DPL achieves the highest scores on all metrics, with an average relative improvement of 4.29% over the best baselines and six metrics showing statistically significant gains. Similar trends are observed on Movies & TV and Books, where DPL also ranks first across all metrics, with 6 and 4 metrics achieving significance, respectively. This result underscores its enhanced personalized text generation capabilities, which can be attributed to its focus on extracting inter-user differences to improve LLM personalization through the selective comparison mechanism and difference extraction standard.
@@ 4.3 Ablation Studies (RQ2)
Random. This variant randomly selects users who have reviews on the same retrieved historical item without considering any specific criteria.
@@
Methods | R-1 | MET. | S-72B |
DPL | |||
Random | |||
SimRank | |||
None | |||
WriOnly | |||
EmoOnly | |||
SemOnly |
Table 2 illustrates the comparison results on ROUGE-1, METEOR, and S-72B, from which we draw the following observations:
@@
For the structured difference extraction, removing parts of the extracted dimensions leads to a general performance drop. This finding suggests that each dimension plays a crucial role in capturing different facets of the differences, thereby supporting meaningful and well-structured extraction for distinguishing task-relevant content.
@@ 4.4.1 Hyper-parameter Analysis
In our investigation, the hyperparameter , which denotes the number of selected representative users for comparison, plays a pivotal role in determining the effectiveness of DPL. We vary within the set to assess the impact of varying these parameters on performance. Additionally, we present the best baseline metrics for comparative analysis, with the results shown in Figure 2, using a fixed temperature of 0.8. Our findings indicate that, generally, increasing the number of representative users leads to slight to moderate improvements in performance, with the best results occurring at . This underscores that effective difference extraction depends on selecting an appropriate number of comparative representatives.
@@ 4.4.2 Impact of User Uniqueness
Following this classification, we compare the performance of DPL with the non-personalization method across the two groups. For simplicity, we report the absolute improvement of ROUGE-1 and S-72B, as the trends for other metrics are consistent. As illustrated in Figure 3, the unique group shows a significantly greater improvement than the non-unique group. This suggests that the more unique the user, the better we can capture insights into their differences relative to others, aligning with the philosophy that “differences make us unique”.
@@ 4.5 Case Study (RQ5)
Finally, we conduct a case study to gain a deeper understanding of DPL’s capacity for personalized review generation. Specifically, we select a representative item from the Movies & TV dataset and compare a single review generated by DPL with that produced by the representative baseline method, LLM-TRSR. The detailed reviews are presented in Figure 4, where we highlight consistent sections between the generated reviews and the user’s real review by distinct colors.
We observe that the review generated by DPL exhibits more consistent sections. Furthermore, it offers a more personalized perspective and concise content, whereas the review generated by LLM-TRSR adopts a more objective stance and contains more tedious information. The alignment of DPL with the user’s writing style and semantic preferences further supports its superiority. More case studies are provided in Appendix H.
@@ 5 Related Work
With the widespread use of LLMs Zhang et al. (2024b); Li et al. (2025); Zhao et al. (2025b); Liu et al. (2025b); Du et al. (2025); Fang et al. (2025), there is a growing need for personalized responses tailored to individual users. LLM personalization has been explored across various domains, underscoring its importance Zhao et al. (2025c, a); Zhang et al. (2025a); Cai et al. (2025). While existing works advance personalization in different scenarios, such as recommendation Chen et al. (2024b); Bao et al. (2024); Zhang et al. (2024a) and image generation Shen et al. (2024); Xu et al. (2025a, c), personalized text generation remains the foundation of LLM personalization Zhang et al. (2024d). LaMP Salemi et al. (2024b) evaluates short-form personalized text, later extended by LongLaMP Kumar et al. (2024) to assess long-form text generation. Among these, personalized item review writing stands out as a key task Peng et al. (2024); Au et al. (2025) since user reviews naturally reflect individual preferences Ni et al. (2019), serving as a critical scenario to evaluate a model’s ability to generate personalized content.
Recent LLM-based methods enhance personalization using user-specific content, which can be categorized into: 1) retrieval-based methods, where ROPG Salemi et al. (2024a) introduces a retrieval selection model to dynamically choose the best retriever and HYDRA Zhuang et al. (2024) employs a personalized reranker to prioritize useful information; and 2) summarization-based methods, where ONCE Liu et al. (2024b) generates user profiles by summarizing topics and regions of interest in their browsing history and PPlug Liu et al. (2024a) converts user history into personalized embeddings.
Besides, a recent study Persona-DB Sun et al. (2025) uses data from other users to enhance personalized text generation. However, this method mainly focuses on enriching relevant knowledge rather than capturing each user’s uniqueness, key to effective personalization. To leverage inter-user data, contrastive learning Gao and Das (2024) offers a promising way of using both positive and negative examples to better describe intents. However, its performance relies on meaningful negative examples, which are challenging to obtain in user-centered LLM personalization. This reliance limits the full leverage of inter-user information for LLM personalization. To the best of our ability, our proposed DPL addresses this limitation.
@@ 6 Conclusions
In our research, we use an open-source dataset, ensuring full compliance with the original dataset’s MIT license. Our research prioritizes data integrity, privacy, and ethical considerations to uphold the responsible and fair use of open-source resources.
@@ Appendix A Dataset Details
Books (839) | Train | 2536 | 29.3422.63 | 1295.77827.48 |
Val | 317 | 33.8422.51 | 1216.49768.44 | |
Test | 317 | 34.8422.51 | 1194.90801.18 |
@@
Data Curation. We perform a detailed data-cleaning process to create a high-quality benchmark. First, we ensure that each review includes a corresponding review_title, rating, and timestamp, with a minimum text length of 200 characters to meet the long-form criteria. For each item, we verify that it has an item_title and that its description falls within the 100 to 2000-character range to maintain high-quality content. Following this, we ensure that each product has at least four unique reviewers and retain only the most recent review from each reviewer for a given item. Additionally, we restrict the number of reviews per reviewer to between 18 and 500 to filter out users who may engage in review manipulation or have insufficient data for meaningful personalization. To meet the requirements of our experimental setup, we further ensure that each user in the validation and test sets has at least 8 historical records, where each record corresponds to an item that has been reviewed by at least 5 users, including the target user. This constraint guarantees that we can retrieve a sufficient number of other users for computing difference-aware representations during evaluation. These settings are designed to align with our experiments.
Dataset Usage. In our experiments, we utilize only the test dataset for a direct evaluation of model performance. The details of model inputs are provided in Appendix G. Notably, the training and validation datasets are included to support related research.
Appendix B Baselines Details